# Inhibition of splitting of the chiral and deconfinement transition due to rotation in QCD: the phase diagram of linear sigma model coupled to Polyakov loop

### v1

@@ Abstract
We discuss the effect of rigid rotation on critical temperatures of deconfinement and chiral transitions in the linear sigma model coupled to quarks and the Polyakov loop. We point out the essential role of the causality condition, which requires that any point of the system should rotate slower than the velocity of light. We show that imposing this physical requirement leads to inhibition of the splitting between the chiral and confining transitions, which becomes negligibly small ($\Delta T\sim 1$ MeV or less) for experimentally relevant, slow angular velocities $\Omega\sim 10$ MeV of a $(5-10)$ fm-sized systems. Moreover, the boundedness of the system has a much bigger effect on temperature splitting than the rotation itself: the splitting reaches 10 MeV in a small, one-fermi-sized non-rotating system. The temperature splitting may, however, become enhanced in an academic limit of ultra-relativistic regimes when the boundary of the system rotates at near-to-light velocities.
@@ I Introduction
The situation on the theoretical side can be improved by phenomenologically tuning the parameters of the system in such a way that the parameters of the original model become dependent on the angular velocity itself Sun et al. (2024). However, similarly to the first-principle numerical approaches, the fine-tuned effective infrared models offer somewhat limited help in elucidating the physical mechanism that lies behind this fine-tuning. On the other hand, the inclusion of couplings that depend on the external parameters (temperature, rotation) aligns well with the logic of the Polyakov-loop improved models in which the Lagrangian that incorporates the dynamics of the loop contains already temperature-dependent couplings Fukushima (2004); Roessner et al. (2007); Farias et al. (2017). In our paper, we consider the traditional sigma model in which the model parameters and couplings do not depend on the external parameters.
@@
Unfortunately, the PLSMq cannot describe the chiral (or chiral and deconfinement) phase transitions even qualitatively. However, this model, similarly to the one considered in Ref. Sun et al. (2023b), serves as a very successful effective model for characterizing the infrared phenomenology of the non-rotating quark-gluon matter. In this paper, we use it to probe the effects associated with the spatial boundedness of the system in the presence of rotation, and confront them with the results in the unbounded case Sun et al. (2023b). Moreover, we expect, given our results below, that the modification of the effective infrared model by introducing environment-dependent couplings will not modify the qualitative conclusions of our paper.
@@ II.1 Lagrangian
The Polyakov-loop enhanced linear sigma model coupled to quarks (PLSMq) has three types of variables: the quark doublet field $\psi(x)$, the $O(4)$ chiral fields $\phi=(\sigma,\vec{\pi})$ with the pion isotriplet $\vec{\pi}=(\pi^{+},\pi^{-},\pi^{0})$, and the Polyakov loop variable $L(x)$. They enter the Lagrangian of the model via the following three terms that are associated with the mentioned variables:
@@ II.1.1 Chiral sector
exhibits both spontaneous and explicit breaking of chiral symmetry. The potential gives a non-zero expectation value for the $\sigma$ field and provides the masses $m_{\sigma}$ and $m_{\pi}$ to the sigma and the pion mesons, respectively. Writing $\sigma=\langle\sigma\rangle+\delta\sigma$ and $\vec{\pi}=\langle\vec{\pi}\rangle+\delta\vec{\pi}$, these masses can be obtained from the expansion in terms of small fluctuations $\delta\sigma$ and $\delta\pi$:
@@
with $V_{0}=V(\langle\sigma\rangle,0)=\frac{\lambda}{4}(\langle\sigma\rangle^{2}-v^{2})^{2}-h\langle\sigma\rangle$ and
@@
where we imposed $\langle\vec{\pi}\rangle=0$ as it follows from the minimum of potential (3).
The analysis of the linear sigma model is usually done in a mean-field approach, in which the “slow” meson sector is considered a classical field. In contrast, quarks are treated as quantum fields that represent “fast” degrees of freedom. The parameters of Lagrangian (2) are chosen to match the low-energy phenomenology Scavenius et al. (2001b). The masses in Eq. (3) serve as the model parameters that coincide with corresponding meson masses in the vacuum, i.e., at zero temperature, $T=0$, and in the absence of rotation, $\Omega=0$. In particular, $m_{\pi}\approx 138\,\mbox{MeV}$, $\langle\sigma\rangle=f_{\pi}\approx 93\,\mbox{MeV}$. In the absence of electromagnetic interactions, the masses of charged, $\pi^{\pm}$, and neutral, $\pi^{0}$, mesons do not split.
@@ II.1.2 Confining sector
$\Phi$ $={\mathcal{P}}\exp\Bigl{[}i\int\limits_{0}^{1/T}{\rm d}\,x_{4}\mathcal{A}_{4}(\vec{x},x_{4})\Bigr{]}\,.$ (6)
The integration takes place along a closed loop in the compactified imaginary time $\tau\equiv x_{4}$, where $\mathcal{A}_{4}=i\mathcal{A}_{0}$ is the matrix-valued temporal component of the Euclidean gauge field $\mathcal{A}_{\mu}$ and the symbol ${\mathcal{P}}$ denotes path ordering that insures the gauge-covariance of the whole expression (6). In the PLSMq, the Polyakov loop $L$ plays the role of a homogeneous classical scalar field, which can be associated with a coordinate-independent gluon field $\mathcal{A}_{4}=t_{3}\,{\mathcal{A}}_{4}^{(3)}+t_{8}\,{\mathcal{A}}_{4}^{(8)}$ as $\Phi=\exp(i\mathcal{A}_{4}/T)$, where $t_{3}$ and $t_{8}$ are the diagonal generators of the $SU(3)$ gauge group.
@@
Definition (6) of the Polyakov loop implies that the value of this quantity is limited to the interval $|L^{*}|\leqslant 1$ and $|L|\leqslant 1$, which is also consistent with a logarithmic divergence of the potential (8). According to the parameter set (12), the value $L^{*}=L=1$ can only be reached in the limit $T\rightarrow\infty$. In the confined phase, $T<T_{0}$, the potential has one trivial minimum, $L=0$. As the temperature rises and the system undergoes the phase transition, the single minimum at $L=0$ of the potential splits into three degenerate minima labeled by the ${\mathbb{Z}}_{3}$ variable. Thus, in the deconfined phase, the ${\mathbb{Z}}_{3}$ symmetry is spontaneously broken, and $\langle L\rangle\neq 0$.
@@ II.2 Thermodynamics of nonrotating system
Before considering the main subject of our paper, the quark system in rotation, we first describe comprehensively a general principle how the thermodynamics of the non-rotating plasma determined in the PLSMq model. A more detailed review and careful determination of the thermodynamical properties of the model will be reported in our forthcoming paper Singha et al. (2024).
@@ II.2.1 Thermodynamic potential
We adopted the mean-field approximation by replacing $\sigma$ and ${\vec{\pi}}$ by their mean-field (expectation) values. As we discussed above, in the ground state, $\langle\vec{\pi}\rangle\equiv 0$ in all studied phases. The quantity $\langle\sigma\rangle$ determines the strength of the chiral symmetry breaking and can be nonzero.
@@
The last term in Eq. (16) corresponds to the quark contribution to the thermodynamic potential, which is given by the last multiplier in Eq. (15). The quark part can be split into its vacuum and thermal constituents: $F_{\psi\bar{\psi}}=F^{\rm vac}_{\psi\bar{\psi}}+F^{\beta}_{\psi\bar{\psi}}$, where the vacuum contribution to the free energy,
@@
is infinite. We, however, follow the arguments provided in Ref. Scavenius et al. (2001b) and ignore below the vacuum quark part by setting $F^{\rm vac}_{\psi\bar{\psi}}=0$. Since now $F_{\psi\bar{\psi}}=F^{\beta}_{\psi\bar{\psi}}$, we will omit the “thermal” $\beta$ superscript and instead identify the fermionic free energy entirely with its thermal contribution Scavenius et al. (2001b); Ratti et al. (2007):
@@
where we introduced the particle and antiparticle free energies, respectively (see Ref. Torres-Rincon and Aichelin (2017) for technical details):
@@ II.2.2 Saddle point approximation
The second restriction in Eq. (28) also has two bounds. The lower bound appears due to the property that in the vacuum, the Polyakov loop expectation value vanishes, $L=L^{*}=0$, reflecting the fact that the addition of an infinitely heavy (anti-)quark requires an infinitely large change in the free energy of the system. Therefore, the corresponding free energies, of quark, $F_{Q}=-T\ln L$, and of an anti-quark, $F_{\bar{Q}}=-T\ln L^{*}$, are infinite. Notice that within the scope of the Polyakov loop models, the two conjugate fields $L$ and $L^{*}$ are treated as real but independent numbers. It means, in particular, that the free energy of a heavy (anti-)quark, defined above, is a real number.
As the medium becomes more energetic (either denser, at higher chemical potential $\mu$; or hotter, at higher temperature $T$), the quark-gluon medium enters a deep deconfinement phase, which is achieved when the Polyakov loop approaches its maximal expectation value of $L,L^{*}\rightarrow 1$. Values outside the intervals indicated in Eq. (28) cannot represent physically realizable systems. The upper bound on the Polyakov loop (28) arises due to the mathematical definition of the Polyakov loop (6).
@@ II.3 Phase diagram of non-rotating plasma
For a static plasma, the thermodynamic limit has a well-defined meaning, which allows us to check our algorithms for determining the phase diagram before proceeding, in the next section, to the more complicated case of rigidly rotating plasma.
@@
Figure 1 illustrates a typical behavior of the full thermodynamic potential (16) – excluding the vacuum part – as a function of the $\sigma$ mean field at the vector chemical potential $\mu=0.25$ GeV, which is sufficiently large to allow for a first-order phase transition. In this figure, we consider the field $\sigma$ as a free (external) parameter and find the values of the $L$ and $L^{*}$ that minimize $F$ at the fixed values of $\sigma$, $T$, and $\mu$. In the vicinity of the phase transition, the free energy $F$ develops multiple extrema corresponding to two minima and one maximum. The thermodynamically favored phase corresponds to the global minimum, while the other minimum represents a metastable state and the maximum gives us an unstable point.
At this particular value of the vector chemical potential, $\mu=0.25$ GeV, the phase transition appears at $T_{c}\simeq 0.15$ GeV, implying that the theory stays in the chirally-broken phase at lower temperatures and in the chirally-restored phase at higher temperatures. Notice that the chiral restoration in the high-temperature phase has an approximate meaning because both these phases are analytically connected. Indeed, Fig. 1 shows that at the phase transition point, the mean value of the field $\sigma$ drops rather strongly but still does not vanish at the high-temperature phase at $T>T_{c}$. The same remark applies also to the deconfinement transition.
@@
Figure 2 shows the order parameter for the chiral transition, $\sigma$, normalized to its vacuum value $\sigma_{\rm vac}=f_{\pi}$ as a function of temperature. In the same figure, we also show the expectation value of the Polyakov loop, which is the confinement order parameter. This figure clearly shows that the increase in temperature or chemical potential drives the system from the confinement (chirally broken) to the deconfinement (chirally restored) phase, as expected.
Notice that the presence of finite density of quark matter, characterized by the non-zero chemical potential, $\mu\neq 0$, leads to the explicit breaking of the charge-conjugation symmetry. In other words, the properties of quarks become different from the properties of anti-quarks. In particular, the free energy of an infinitely heavy quark, $F_{Q}=-T\ln L$, and the free energy of an anti-quark, $F_{\bar{Q}}=-T\ln L^{*}$, differ from each other (we remind that both $L$ and $L^{*}$ are real-valued quantities). Figure 2 illustrates this behavior: adding another quark to a quark-rich medium (with $\mu>0$) is more costly than adding an anti-quark to the same medium (in other words, $F_{Q}>F_{\bar{Q}}$ for all values of $\mu>0$).
@@
In the absence of a discontinuity at low $\mu$’s, the system goes continuously from one phase to the other as parameters change. In practical terms, discrimination between a first-order transition and a sharp crossover transition poses a formidable numerical challenge. Pinpointing the exact position of the pseudo-critical parameters of the crossover transition has a certain degree of arbitrariness, which depends on the choice of the observables used to identify the crossover.
@@
In our work, the point of the crossover is identified as an inflection point with respect to the relevant order parameter $\sigma$, $L$, and $L^{*}$. The inflection is searched along a straight line that is connected with the origin of the parameter space, $\mu=T=0$. The resulting phase diagram is shown in Fig. 3. Qualitatively, it agrees with known results obtained in the thermodynamic limit Schaefer et al. (2007). [^1] The predictions of the effective models may differ quantitatively depending on number quark flavors, values of the model couplings, treatment of field fluctuations and implementation of renormalization, (see, for example, Ref. Kashiwa et al. (2008)). There is no splitting between the deconfining and chirality-restoring transitions neither for thermodynamic phase transition at higher $\mu$ nor for the crossover regime at the lower-density part of the phase diagram at smaller values of the chemical potential $\mu$. As we will see below, this property will be lost for a system where two of the three spatial directions are bounded, as it is pertinent to the rotating systems.

### v2

@@ Abstract
We discuss the effect of rigid rotation on the critical temperatures of deconfinement and chiral transitions in the linear sigma model coupled to quarks and the Polyakov loop. We point out the essential role of the causality condition, which requires that any point of the system should rotate slower than the velocity of light. We show that imposing this physical requirement leads to inhibition of the splitting between the chiral and confining transitions, which becomes negligibly small ($\Delta T\sim 1$ MeV or less) for experimentally relevant, slow angular velocities $\Omega\sim 10$ MeV of a $(5-10)$ fm-sized systems. Moreover, the boundedness of the system has a much bigger effect on temperature splitting than the rotation itself: the splitting reaches 10 MeV in a small, one-fermi-sized non-rotating system. The temperature splitting may, however, become enhanced in an academic limit of ultra-relativistic regimes when the boundary of the system rotates at near-to-light velocities.
@@ I Introduction
The situation on the theoretical side can be improved by phenomenologically tuning the parameters of the system in such a way that the parameters of the original model become dependent on the angular velocity itself Sun et al. (2024). However, similarly to the first-principle numerical approaches, the fine-tuned effective infrared models offer somewhat limited help in elucidating the physical mechanism that lies behind this fine-tuning. On the other hand, the inclusion of couplings that depend on the external parameters (temperature, rotation) aligns well with the logic of the Polyakov-loop improved models, in which the Lagrangian that incorporates the dynamics of the loop contains already temperature-dependent couplings Fukushima (2004); Roessner et al. (2007); Farias et al. (2017).
One should also comment on the potential risk of double counting of the gluonic degrees of freedom when incorporating the Polyakov loop terms into effective quark models, such as the Nambu–Jona-Lasinio Nambu and Jona-Lasinio (1961a, b) or quark-meson sigma models. For example, in the linear sigma model with quarks Gell-Mann and Lévy (1960), formulated at zero temperature, the terms that provide the interaction of mesons with quarks and the self-interaction of the mesons are interpreted as a result of the dynamics of gluons. In other words, in this and similar models, the gluon degrees of freedom are already considered to be integrated out. Therefore, an additional interaction of the quark fields with the gluon field corresponding to the Polyakov loop may be regarded as a redundancy. However, one can justify the presence of these terms by arguing that the original quark-meson model Gell-Mann and Lévy (1960) has been developed to describe zero-temperature physics, which does not account for the global center symmetry ${\mathbb{Z}}_{3}$. The breaking of this symmetry is associated with the finite-temperature deconfinement transition. The information about this transition is carried by the Polyakov loop, which appears only in a finite-temperature system. Moreover, the time-like gauge field should be interpreted as a non-trivial classical background, capturing essential features of the confinement-deconfinement transition in QCD. In contrast, the residual gluon fluctuations, which are integrated out in this and similar models, are accounted for through their contribution to the effective four-fermion interaction terms or the interaction with colorless bound states, such as light mesons. Therefore, the inclusion of the Polyakov loop in the finite-temperature effective model seems to be a legitimate extension of the zero-temperature version of the model Fukushima and Skokov (2017).
In our paper, we consider the traditional sigma model in which the model parameters and couplings do not depend on external parameters.
@@
Unfortunately, the PLSMq cannot describe the chiral (or chiral and deconfinement) phase transitions even qualitatively. However, this model, similarly to the one considered in Ref. Sun et al. (2023b), serves as a very successful effective model for characterizing the infrared phenomenology of the non-rotating quark-gluon matter. In this paper, we use it to probe the effects associated with the spatial boundedness of the system in the presence of rotation and confront them with the results in the unbounded case Sun et al. (2023b). Moreover, we expect, given our results below, that the modification of the effective infrared model by introducing environment-dependent couplings will not modify the qualitative conclusions of our paper.
@@ II.1 Lagrangian
The Polyakov-loop enhanced linear sigma model coupled to quarks (PLSMq) has three types of variables: the quark spinor $\psi(x)$ accounting for $N_{f}=2$ flavours and $N_{c}=3$ colours, the $O(4)$ chiral fields $\phi=(\sigma,\vec{\pi})$ with the pion isotriplet $\vec{\pi}=(\pi^{+},\pi^{-},\pi^{0})$, and the Polyakov loop variable $L(x)$. They enter the Lagrangian of the model via the following three terms that are associated with the mentioned variables:
@@ II.1.1 Chiral sector
exhibits both spontaneous and explicit breaking of chiral symmetry. The potential gives a non-zero expectation value $\langle\sigma\rangle=f_{\pi}$ for the $\sigma$ field and a vanishing pion expectation value, $\langle\vec{\pi}\rangle=0$, providing the masses $m_{\sigma}$ and $m_{\pi}$ to the sigma and the pion mesons. Writing $\sigma=\langle\sigma\rangle+\delta\sigma$ and $\vec{\pi}=\langle\vec{\pi}\rangle+\delta\vec{\pi}$, these masses can be obtained from the expansion in terms of small fluctuations $\delta\sigma$ and $\delta\pi$:
@@
with $V_{0}=V(\langle\sigma\rangle,\langle\vec{\pi}\rangle)=\frac{\lambda}{4}(\langle\sigma\rangle^{2}-v^{2})^{2}-h\langle\sigma\rangle$ and
@@
where we imposed $\langle\vec{\pi}\rangle=0$ as it follows from the minimum of the potential (3).
The analysis of the linear sigma model is usually done in a mean-field approach, in which the “slow” meson sector is considered a classical field. In contrast, quarks are treated as quantum fields that represent “fast” degrees of freedom. The parameters of the Lagrangian (2) are chosen to match the low-energy phenomenology Scavenius et al. (2001b). The masses in Eq. (5) serve as the model parameters that coincide with corresponding meson masses in the vacuum, i.e., at zero temperature, $T=0$, and in the absence of rotation, $\Omega=0$: $m_{\pi}=138\,\mbox{MeV}$ and $m_{\sigma}=600\,\mbox{MeV}$. In the absence of electromagnetic interactions, the masses of the charged, $\pi^{\pm}$, and neutral, $\pi^{0}$, mesons do not split.
@@ II.1.2 Confining sector
$\Phi$ $={\mathcal{P}}\exp\Bigl{[}-i\int\limits_{0}^{1/T}{\rm d}\,x_{4}\mathcal{A}_{4}(\vec{x},x_{4})\Bigr{]}\,.$ (6)
The integration takes place along a closed loop in the compactified imaginary time $\tau\equiv x_{4}$, where $\mathcal{A}_{4}=i\mathcal{A}_{0}$ is the matrix-valued temporal component of the Euclidean gauge field $\mathcal{A}_{\mu}$ and the symbol ${\mathcal{P}}$ denotes path ordering that insures the gauge-covariance of the whole expression (6). In the PLSMq, the Polyakov loop $L$ plays the role of a homogeneous classical scalar field, which can be associated with a coordinate-independent gluon field $\mathcal{A}_{4}=\frac{1}{2}(t_{3}\,{\mathcal{A}}_{4}^{(3)}+t_{8}\,{\mathcal{A}}_{4}^{(8)})$ as $\Phi=\exp(-i\mathcal{A}_{4}/T)$, where $t_{3}$ and $t_{8}$ are the diagonal generators of the $SU(3)$ gauge group.
@@
The definition (6) of the Polyakov loop implies that the value of this quantity is limited to the interval $|L^{*}|\leqslant 1$ and $|L|\leqslant 1$, which is also consistent with a logarithmic divergence of the potential (8). According to the parameter set (12), the value $L^{*}=L=1$ can only be reached in the limit $T\rightarrow\infty$. In the confined phase, $T<T_{0}$, the potential has one trivial minimum, $L=0$. As the temperature rises and the system undergoes the phase transition, the single minimum at $L=0$ of the potential splits into three degenerate minima labeled by the ${\mathbb{Z}}_{3}$ variable. Thus, in the deconfined phase, the ${\mathbb{Z}}_{3}$ symmetry is spontaneously broken, and $\langle L\rangle\neq 0$.
One can argue that in the presence of vorticity, the rotation effects should be incorporated into the Polyakov loop potential because this potential describes the (de)confinement phenomenon in SU(3) gauge theory while, at the same time, the deconfinement transition is affected by the rotation of the same pure gauge matter Braguta et al. (2020, 2021). This modification of the potential is related to the very fundamental non-perturbative mechanism of confinement and, therefore, cannot be taken into account analytically. [^1] For a perturbative calculation of a rotational contribution to an effective potential of the Polyakov loop, see Refs. Chen et al. (2022, 2024). Hence, the rotation-dependent terms in the potential should be introduced in a phenomenological way, as it was done in Ref. Sun et al. (2023a). By promoting the parameters of the Polyakov loop potential to phenomenological functions of the angular velocity, one finds that the model, indeed, better describes a particular, volume-averaged set of available lattice data for vortical gluon matter.
However, the recent lattice data have demonstrated that the rotation generates inhomogeneities of the Polyakov loop expectation values in the radial direction with respect to the axis of rotation  Braguta et al. (2024b). This effect cannot be described by translation-invariant terms in the potential that were incorporated in the study of Ref. Sun et al. (2023a). Therefore, in our paper, we follow the standard approach in which the potential does not include additional rotation-dependent contributions (8). This approach will allow us to test how well the traditional model describes the rotational phenomena in the quark-gluon plasma.
@@ II.1.3 Quark sector
Summarizing the description of the system, we discussed an effective infrared model that describes the low-energy dynamics of quark degrees of freedom that are also coupled to the confining gluonic degrees of freedom. The latter is represented by the gauge field ${\mathcal{A}}_{4}$, which is directly related to the confining order parameter, the Polyakov loop. One could think that in this approach, the additional explicit quark-gluon interactions, relevant at $T\neq 0$, might represent an additional redundancy since the gluonic fields are already incorporated in the $T=0$ parameters of the quark-meson model. However, in the adopted approach, these terms are treated differently: the Polyakov loop in quark-gluon terms accounts for the central symmetry that distinguishes the confining and deconfining phases and represents a non-trivial classical gluon background, while the quark-meson and meson-meson terms originate from the $T=0$ gluon fluctuations that are not treated as classical fields.
@@ II.2 Thermodynamics of nonrotating system
Before considering the main subject of our paper, the quark system in rotation, we first comprehensively describe the general principle of how the thermodynamics of the non-rotating plasma is determined in the PLSMq model. A more detailed review and careful determination of the thermodynamic properties of the model will be reported in our forthcoming paper Singha et al. (2024).
@@ II.2.1 Thermodynamic potential
We adopted the mean-field approximation by replacing $\sigma$ and ${\vec{\pi}}$ by their mean-field (expectation) values as we discussed above, in the ground state, $\langle\vec{\pi}\rangle\equiv 0$ in all studied phases. The quantity $\langle\sigma\rangle$ determines the strength of the chiral symmetry breaking and can be nonzero.
@@
The last term in Eq. (16) corresponds to the quark contribution to the thermodynamic potential, which is given by the last multiplier in Eq. (15). The quark part can be split into its vacuum and thermal constituents: $F_{\psi\bar{\psi}}=F^{\rm vac}_{\psi\bar{\psi}}+F^{\beta}_{\psi\bar{\psi}}$, where the vacuum contribution to the thermodynamic potential,
@@
is infinite. We, however, follow the arguments provided in Ref. Scavenius et al. (2001b) and ignore below the vacuum quark part by setting $F^{\rm vac}_{\psi\bar{\psi}}=0$. Since now $F_{\psi\bar{\psi}}=F^{\beta}_{\psi\bar{\psi}}$, we will omit the “thermal” $\beta$ superscript and instead identify the fermionic grand potential entirely with its thermal contribution Scavenius et al. (2001b); Ratti et al. (2007):
@@
where we introduced the particle and antiparticle contributions, respectively (see Ref. Torres-Rincon and Aichelin (2017) for technical details):
@@ II.2.2 Saddle point approximation
The second restriction in Eq. (28) also has two bounds. The lower bound appears due to the property that in the vacuum, the Polyakov loop expectation value vanishes, $L=L^{*}=0$, reflecting the fact that the addition of an infinitely heavy (anti-)quark requires an infinitely large change in the thermodynamic potential of the system. Therefore, the corresponding contributions of one quark, $F_{Q}=-T\ln L$, and of one anti-quark, $F_{\bar{Q}}=-T\ln L^{*}$, are infinite. Notice that within the scope of the Polyakov loop models, the two conjugate fields $L$ and $L^{*}$ are treated as real but independent numbers. It means, in particular, that the free energy of a heavy (anti-)quark, defined above, is a real number.
As the medium becomes more energetic (either denser, at higher chemical potential $\mu$; or hotter, at higher temperature $T$), the quark-gluon medium enters a deep deconfinement phase, which is achieved when the Polyakov loop and its conjugate approach their maximal expectation values, $L,L^{*}\rightarrow 1$. Values outside the intervals indicated in Eq. (28) cannot represent physically realizable systems. The upper bound on the Polyakov loop (28) arises due to the mathematical definition of the Polyakov loop (6).
@@ II.3 Phase diagram of non-rotating plasma
For a static plasma, the thermodynamic limit has a well-defined meaning, which allows us to check our algorithms for determining the phase diagram before proceeding, in the next section, to the more complicated case of the rigidly-rotating plasma.
@@
Figure 1 illustrates the typical behavior of the full thermodynamic potential (16) – excluding the vacuum part – as a function of the $\sigma$ mean field at the vector chemical potential $\mu=0.25$ GeV, which is sufficiently large to allow for a first-order phase transition. In this figure, we consider the field $\sigma$ as a free (external) parameter and find the values of the $L$ and $L^{*}$ that minimize $F$ at fixed values of $\sigma$, $T$, and $\mu$. In the vicinity of the phase transition, the free energy $F$ develops multiple extrema corresponding to two minima and one maximum. The thermodynamically favored phase corresponds to the global minimum, while the other minimum represents a metastable state and the maximum gives us an unstable point.
At this particular value of the vector chemical potential, $\mu=0.25$ GeV, the phase transition appears at $T_{c}\simeq 0.15$ GeV, implying that the theory stays in the chirally-broken phase at lower temperatures and in the chirally-restored phase at higher temperatures. Notice that the chiral restoration in the high-temperature phase has an approximate meaning. Indeed, Fig. 1 shows that at the phase transition point, the mean value of the field $\sigma$ drops rather strongly but still does not vanish at the high-temperature phase at $T>T_{c}$. The same remark applies also to the deconfinement transition.
@@
Figure 2 shows the order parameter for the chiral transition, $\sigma$, normalized to its vacuum value $\sigma_{\rm vac}=f_{\pi}$ as a function of temperature. In the same figure, we also show the expectation value of the Polyakov loop $L$ and its conjugate $L^{*}$, which are the confinement order parameters. This figure clearly shows that the increase in temperature or chemical potential drives the system from the confinement (chirally broken) to the deconfinement (chirally restored) phase, as expected.
Notice that the presence of finite density of quark matter, characterized by the non-zero chemical potential, $\mu\neq 0$, leads to the explicit breaking of the charge-conjugation symmetry, such that the properties of quarks become different from the properties of anti-quarks. In particular, the free energy of an infinitely heavy quark, $F_{Q}=-T\ln L$, and the free energy of an anti-quark, $F_{\bar{Q}}=-T\ln L^{*}$, differ from each other (we remind that both $L$ and $L^{*}$ are real-valued quantities). Figure 2 illustrates this behavior: adding another quark to a quark-rich medium (with $\mu>0$) is more costly than adding an anti-quark to the same medium (i.e., $F_{Q}>F_{\bar{Q}}$ for all values of $\mu>0$).
@@
In the absence of a discontinuity at low $\mu$’s, the system goes continuously from one phase to the other as the thermodynamic parameters change. In practical terms, discriminating between a first-order transition and a sharp crossover transition poses a formidable numerical challenge. Pinpointing the exact position of the pseudo-critical parameters of the crossover transition has a certain degree of arbitrariness, which depends on the choice of the observables used to identify the crossover.
@@
In our work, the point of the crossover is identified as an inflection point with respect to the relevant order parameters $\sigma$, $L$, and $L^{*}$. The inflection is searched along a straight line that is connected with the origin of the parameter space, $\mu=T=0$. The resulting phase diagram is shown in Fig. 3. Qualitatively, it agrees with known results obtained in the thermodynamic limit Schaefer et al. (2007). [^2] The predictions of the effective models may differ quantitatively depending on number of quark flavors, values of the model couplings, treatment of field fluctuations and implementation of renormalization (see, for example, Ref. Kashiwa et al. (2008)). There is no splitting between the deconfining and chirality-restoring transitions neither for the thermodynamic phase transition at higher $\mu$ nor for the crossover regime at the lower-density part of the phase diagram at smaller values of the chemical potential $\mu$. As we will see below, this property will be lost for a system where two of the three spatial directions are bounded, as it is pertinent to the rotating systems.
At the low-$T$ high-$\mu$ corner of the phase diagram of Fig. 3, a strong chiral transition, which is marked by a substantial discontinuity in the chiral condensate, is accompanied by very modest discontinuities in the Polyakov loop $L$ as well as its conjugate $L^{*}$. Therefore, the free energies of a heavy quark and a heavy anti-quark experience a relatively insignificant change across the transition. This observation provokes the natural question of whether this transition can be interpreted as a deconfining transition. This question is more than legitimate as the Polyakov loop, strictly speaking, does not play the role of an order parameter in the presence of dynamical quarks that break the center ${\mathbb{Z}}_{3}$ symmetry explicitly. Physically, it means that even at vanishing chemical potential, the dynamical quark fields screen the chromoelectric field of a heavy test quark $Q$. The process proceeds via the breaking of the confining string attached to the heavy quark through the $q\bar{q}$ pair creation. Therefore, the free energy associated with the Polyakov loop $L$ corresponds to the free energy of the $Q\bar{q}$ heavy-light meson, which takes a finite value even in the hadronic phase. Consequently, the expectation value of the Polyakov loop, strictly speaking, does not vanish in the hadron phase Greensite (2003).
This effect becomes even more pronounced in the presence of a finite chemical potential, as the Polyakov loop gets screened by the on-shell quarks that are already present in the system. Moreover, when the density of quarks is high, the mean inter-quark density is smaller than the distance at which the QCD string gets formed. Therefore, the confining property is, in a rigorous sense, lost as the confining string of the heavy quark cannot develop in the system. This analogy with QCD has a qualitative nature since the effective PLSMq model does not possess the dynamical gluon degrees of freedom responsible for the string formation. Still, using continuity arguments that connect analytically the low-$T$ high-$\mu$ corner of the phase diagram with the high-temperature plasma, we interpret the weak discontinuity in the Polyakov loop as a signature of a deconfining transition. It is necessary to mention that at the high quark density, the phase structure of QCD is expected to become more complex as the ground state may be described by more involved quark condensates Fukushima and Sasaki (2013). We do not expand further our speculations about the high-density and low-temperature corner of the QCD phase diagram as this parameter goes beyond the applicability of the simple low-energy models such as PLSMq.